A wood-burning stove reaches its constant surface temperature of 198°C, and the room temperature warms up to a constant temperature of 29°C. The stove has an emissivity of 0.90 and a surface area of 3.50 m2.

a) Determine the radiant power emitted by the stove. b) Determine the net radiant power of the wood stove.

4.

a) Determine the radiant power emitted by the stove. b) Determine the net radiant power of the wood stove.
Answer #1

Stefan-boltzmann constant = = 5.67 x 10-8 W/(m2.K4)

Temperature of the wood stove = T1 = 198 oC = 198 + 273 K = 471 K

Temperature of the room = T2 = 29 oC = 29 + 273 K = 302 K

Surface area of the wood stove = A = 3.5 m2

Emissivity of the wood stove = e = 0.9

Radiant power emitted by the stove = Q

Q = eAT14

Q = (5.67x10-8)(0.9)(3.5)(471)4

Q = 8789.76 W

Net radiant power of the stove = Qnet

Qnet = eA(T14 - T24)

Qnet = (5.67x10-8)(0.9)(3.5)(4714 - 3024)

Qnet = 7304.1 W

a) Radiant power emitted by the stove = 8789.76 W

b) Net radiant power of the wood stove = 7304.1 W